# Family Matters: Language Transfer and Merging for Adapting Small LLMs to Faroese

### v1

@@ 3.2 Merging
Merging provides an efficient alternative to multi-task training by combining different fine-tuned models into one checkpoint, often improving generalization (Wortsman et al., 2022; Yadav et al., 2024). Task Arithmetic merging (Ilharco et al., 2023) computes task vectors as the difference between a fine-tuned model and its base, averages them, and adds the result back to the base model. TIES (Yadav et al., 2023) refines this by keeping only the most influential parameters, resolving sign conflicts, and using a disjoint mean to reduce interference. We apply TIES to languages instead of tasks, using Mergekit (Goddard et al., 2024): starting from SmolLM, we fine-tune on each language to obtain language vectors, merge them, and train the resulting model on Faroese. Since we continue pre-training on Faroese after merging and have limited resources, we select three promising merges that cover different language mixes and weightings: Mergeeq where we merge all five models equally (with pre-normalization weight 1 and density 0.5), Mergeis+ where we merge with bias towards Icelandic (weight 1 for Icelandic, 0.5 for all others, density 0.5), and Merge2, where we merge only two models: Icelandic and Danish (both with weight 1 and density 0.5), as Danish is the mainland Scandinavian language with the lowest perplexity after Faroese continued pre-training (see Table 3(b)).
@@ 3.3 Automatic Evaluation
We evaluate the perplexity on the validation set of the Faroese portion of Fineweb-2 both zero-shot and after continued pre-training on Faroese. In addition, we introduce two benchmarks: FoBLiMP for linguistic acceptability probes, and FoBCoMP for text comprehension probes. We report results on the original SmolLM models as a baseline.
@@
To probe zero-shot linguistic skills, we use minimal pairs with one correct and one corrupted sentence, measuring the percentage of times the model assigns higher probability to the correct sentence. This collection is called FoBLiMP (Faroese Benchmark of Linguistic Minimal Pairs).
To evaluate subject-verb agreement, we use the Faroese portion of MultiBLiMP (Jumelet et al., 2025), containing 232 sentences. ScaLA (Nielsen, 2023) contains sentences corrupted by swapping or deleting words. Originally a binary classification task, we convert it to minimal pairs by realigning correct and incorrect sentences using Levenshtein distance ($\geq 0.85$), with unmatched samples added manually. Concatenating all subsets gives 552 pairs for flip_neighbours and 601 pairs for delete. GermDetect Michael and Horbach (2025) provide automatically corrupted sentences with verb placement errors. After removing pairs with no corruption, we obtain 2,026 pairs. As Faroese allows flexible word order, some corruptions are grammatical, but we conclude from an inspection that the original sentence is mostly more common. We also construct minimal pairs from a human evaluation in Scalvini et al. (2025a), where two raters annotated errors in English-to-Faroese translations from four models. We pair translations with an error difference of at least 2, keeping those with no more than four errors in the better translation and excluding translations containing foreign scripts. This yields 680 pairs.
@@
Evaluating small LLMs in text comprehension is particularly challenging because evaluations often mix formal competence (e.g., grammar) with functional competence (e.g., following prompts) (Kydlíček et al., 2024). Limited fine-tuning data further complicates comparisons. To address this, we also use text comprehension probes in a minimal-pair format. We introduce a set of five probes, called FoBCoMP (Faroese Benchmark of Text Comprehension Minimal Pairs).
We adapt the Faroese news sentiment dataset (Debess et al., 2024) (original labels: positive, negative, neutral) into minimal pairs by adding a sentiment-bearing sentence (“Hetta er gott/ringt”). Neutral labels are excluded as initial experiments showed that words such as “neutral” are never the most probable choice. We evaluate sentence- and article-level samples, keeping only items annotators of the original dataset agreed on, resulting in 91 sentence-level (55 positive, 36 negative) and 84 article-level (51 positive, 33 negative) pairs. Using the same dataset, we filter GPT-4-assigned topic labels confirmed by a human. Minimal pairs consist of one correct topic and one incorrect topic (not assigned to the article), with related-topic pairs curated to make the task realistic (e.g., Local News vs. International News). This yields 234 topic classification pairs. We also adapt the extractive QA dataset FoQA (Simonsen et al., 2025) (2,000 Faroese question–context–answer triplets) into minimal pairs via two methods: (1) Dataset Shuffling: Replacing the correct answer passage with an incorrect but plausible passage from another sample within the context, matching token length. This creates 21,867 pairs. (2) GPT-4 Adversarial Answers: Generating one alternative incorrect answer per sample that is also a span in the dataset, matching token length when possible. Exact length matches occurred for 611 answers; deviations were an average of 1.69 tokens longer than the correct answers. This yields 2,000 pairs.
@@ 3.4 Human Evaluation
As we do not have evaluation sets to assess fine-grained properties for Faroese generation, we perform a human evaluation to assess output quality. The latter requires subjective judgment of linguistic quality, naturalness, and contextual appropriateness. As human evaluation is time-intensive, we focus on four 360M models, which perform among the best. These models vary along two dimensions: full tuning versus LoRA, and transfer from Icelandic versus Mergeis+. The evaluators are two native Faroese speakers, both trained linguists with extensive experience in NLP model evaluation.
@@
Given that we are working with small base models, we chose a simple generation task: sentence continuation. Models are prompted with sentences from a small manually compiled corpus derived from academic papers and local news (not in FineWeb2), with the last words removed and a trailing space. All models produce running text, enabling comparison of output quality. We evaluate outputs for linguistic quality across four subdimensions, scored 0–5: Lexical correctness (valid Faroese words and avoidance of lexical hallucinations), grammatical accuracy (morphological and syntactic correctness, including spelling and typography), Semantic coherence (meaningful, logically consistent content) and fluency/naturalness (native-like expression). The annotators each evaluated 400 continuations (100 prompts, 4 models) We apply a token cut-off of 100 and no penalty for incomplete last sentences.
@@
As zero-shot summarization did not result in reasonable summaries, we create a synthetic dataset for fine-tuning and manually evaluate the summaries generated by the models. The synthetic dataset is compiled with 150 authentic texts in the same domains as the evaluation set (academic, news, blog) paired with summaries generated by Claude Sonnet 4222For fine-tuning details, see Table 7 in Appendix A.. For evaluation, we selected 50 source texts (not part of training text) and generated summaries from all four models, resulting in 200 summary-source pairs. Two annotators conducted a blind evaluation using two criteria: task completion (to what extent did the model solve the task) and linguistic quality, each on a 0–5 scale. We apply a token cut-off of 400 and no penalty for incomplete final sentences.
@@ 4.1 Benchmarks
We first give an overview of the benchmarks results with respect to their difficulty and reliability. FoBLiMP results are shown in Table 1. Models perform well on most linguistic acceptability probes, suggesting these tasks are relatively easy. The main exception is Translation Pairs, where scores are lower, likely due to noise as translation error counts do not always reflect linguistic quality: an error can reflect unrelated aspects such as incorrect content compared to the source sentence.
Full LoRA 135M 360M 135M 360M En 95.25 96.55 96.12 96.55 +Da 96.98 95.25 98.27 96.12 +Is 97.41 99.13 97.41 98.27 +NoB 96.12 100.0 96.98 96.98 +NoN 96.12 98.27 95.68 97.41 +Sv 96.12 98.27 96.98 96.98 Mergeeq 97.41 98.27 96.55 95.25 Mergeis+ 97.84 97.84 97.41 96.55 Merge2 97.84 97.41 97.41 99.13
@@
Full LoRA 135M 360M 135M 360M En 93.48 95.36 92.54 94.66 +Da 94.27 95.11 94.91 95.75 +Is 94.27 95.16 94.57 95.75 +NoB 95.06 95.31 94.91 95.55 +NoN 93.97 94.91 94.47 95.01 +Sv 94.61 95.85 95.36 95.80 Mergeeq 94.66 95.80 95.16 95.31 Mergeis+ 94.57 95.60 95.06 95.26 Merge2 94.52 95.85 94.37 95.60
@@
(f) Mean of the individual FoBLiMP scores in 1(a)-1(e).
@@
Full LoRA 135M 360M 135M 360M En 60.43 68.13 60.43 72.52 +Da 72.52 75.82 61.53 76.92 +Is 63.73 75.82 60.43 69.23 +NoB 75.82 71.42 60.43 72.52 +NoN 76.92 74.72 67.03 71.42 +Sv 70.32 68.13 60.43 69.23 Mergeeq 75.82 72.52 60.43 62.63 Mergeis+ 68.13 80.21 60.43 63.73 Merge2 76.92 75.82 62.63 65.93
@@
Full LoRA 135M 360M 135M 360M En 75.35 84.37 69.57 86.40 +Da 72.86 86.74 69.13 82.93 +Is 75.64 85.46 71.80 86.87 +NoB 72.04 84.16 70.00 86.60 +NoN 72.89 85.05 74.41 86.12 +Sv 70.82 85.12 70.42 86.10 Mergeeq 63.47 83.87 67.43 78.08 Mergeis+ 70.06 83.33 70.63 85.77 Merge2 72.25 85.14 70.19 82.43
@@
(f) Mean of the individual FoBCoMP scores in 2(a)-2(e).
@@
Full LoRA 135M 360M 135M 360M En 73.61 58.27 73.61 58.27 +Da 50.98 40.45 63.23 44.38 +Is 38.77 30.09 40.05 30.54 +NoB 48.59 44.81 56.25 44.51 +NoN 61.58 50.28 68.02 56.62 +Sv 69.96 57.02 78.22 61.43 Mergeeq 103.45 42.40 182.21 338.30 Mergeis+ 94.95 40.16 68.81 62.88 Merge2 65.28 40.39 78.46 146.20
(a) Before continuing training on Faroese (zero-shot).
Full LoRA 135M 360M 135M 360M 4.98 3.75 5.51 4.48 4.19 3.56 4.25 3.55 4.44 3.48 4.53 3.53 4.22 3.63 4.26 3.56 4.60 3.66 4.90 4.08 4.26 3.60 4.21 3.58 4.08 3.41 4.61 3.93 4.08 3.41 4.58 3.77 4.22 3.49 4.56 3.80 arX
(b) After continuing training on Faroese.
Table 3: Average per-token perplexity on the Fineweb-2 evaluation set.
@@ 4.2 RQ1: Transfer Languages are Important
Models adapted via a Scandinavian transfer language consistently show lower perplexities than those adapted directly from English. The difference is larger for smaller models and for LoRA models, suggesting that LoRA benefits more from better-initialized parameters, consistent with Biderman et al. (2024) who find full tuning more sample-efficient than LoRA in domain adaptation.
Table 3(a) shows zero-shot perplexities before Faroese adaptation. Icelandic performs best, followed by Danish and Norwegian-Bokmål; Swedish is barely better than English-only. This likely reflects script overlap: Faroese shares most letters with Icelandic, some diacritics with Danish and Norwegian, but none beyond the English alphabet with Swedish. After Faroese adaptation (Table 3(b)), Icelandic remains best for larger models, while for smaller models, higher-resource mainland Scandinavian languages outperform Icelandic (except for Norwegian-Nynorsk, the lowest-resorce language).
@@
In general,
Icelandic often provides the strongest transfer. It gives the best perplexity before Faroese adaptation and clear advantages in FoBLiMP probes like subject-verb agreement, supporting our choice of Icelandic for human evaluation. Figure 1 shows that choosing Icelandic over English gives almost consistent improvements for FoBLiMP and in many cases gains for FoBCoMP.
Figure 1: Win rates of Icelandic over English only. FoBLiMP (left), FoBCoMP (right).
Figure 2 illustrates how often all transfer setups outperform English, indicating that any additional transfer language is better than none. For FoBLiMP, especially in 360M models, transfer is crucial, while for FoBCoMP, the best language choice is less clear.
Figure 2: Win rates of all transfer setups over English only (i.e., any Scandinavian language is better than none. FoBLiMP (left), FoBCoMP (right).
Comparing single languages for the transfer (Figure 3), the choice is less clear: Icelandic has 14 wins and Danish 13. For FoBLiMP, Icelandic leads with 9 wins versus 7, and 4 ties; for FoBCoMP, Danish leads with 9 wins versus 6, and 5 ties.
Figure 3: Win rates of transfer languages across models and adaptation setups. In case of a tie, we count both. FoBLiMP (left), FoBCoMP (right).

### v2

@@ 3.2 Merging
We apply TIES merging to the continued pre-trained models in different languages, using Mergekit (Goddard et al., 2024): starting from SmolLM, we fine-tune on each language to obtain language vectors, merge them, and train the resulting model on Faroese. We select three promising merges that cover different language mixes and weightings: Mergeeq where we merge all five models equally (with pre-normalization weight 1 and density 0.5), [^3] See Section 2.3 for definitions of merging parameters. resulting in a strong bias towards the mainland Scandinavian languages as four of them are included in the merge; Mergeis+ where we merge with bias towards Icelandic (weight 1 for Icelandic, 0.5 for all others, density 0.5), and Mergeda+is, where we merge only two models: Icelandic and Danish (both with weight 1 and density 0.5), as Danish is the mainland Scandinavian language with the lowest perplexity after Faroese continued pre-training (see Table 5).
@@ 3.3 Automatic Evaluation
We evaluate the perplexity on the validation set of the Faroese portion of Fineweb-2. In addition, we introduce two benchmarks: FoBLiMP for linguistic acceptability probes, and FoBCoMP for text comprehension probes. We report results on the original SmolLM models as a baseline.
@@
To probe zero-shot linguistic skills, we use minimal pairs with one correct and one corrupted sentence, measuring the percentage of times the model assigns higher probability to the correct sentence. This collection is called FoBLiMP (Faroese Benchmark of Linguistic Minimal Pairs). Table 9 in Appendix B provides an overview of sources, subsets and modifications, and statistics. To evaluate subject-verb agreement, we use the Faroese portion of MultiBLiMP (Jumelet et al., 2025), containing 232 sentences. ScaLA (Nielsen, 2023) contains sentences corrupted by swapping or deleting words. Originally a binary classification task, we convert it to minimal pairs by realigning correct and incorrect sentences using Levenshtein distance ($\geq 0.85$), with unmatched samples added manually. Concatenating all subsets gives 552 pairs for flip_neighbours and 601 pairs for delete. GermDetect (Michael and Horbach, 2025) provide automatically corrupted sentences with verb placement errors. After removing pairs with no corruption, we obtain 2,026 pairs. As Faroese allows flexible word order, some corruptions are grammatical, but we conclude from an inspection that the original sentence is mostly more common. We also construct minimal pairs from a human evaluation in Scalvini et al. (2025a), where two raters annotated errors in English-to-Faroese translations from four models. We pair translations with an error difference of at least 2, keeping those with no more than four errors in the better translation and excluding translations containing foreign scripts. This yields 680 pairs.
@@
Evaluating small LLMs in text comprehension is particularly challenging because evaluations often mix formal competence (e.g., grammar) with functional competence (e.g., following prompts) (Kydlíček et al., 2024). Limited fine-tuning data further complicates comparisons. To address this, we also use text comprehension probes in a minimal-pair format. We introduce a set of five probes, called FoBCoMP (Faroese Benchmark of Text Comprehension Minimal Pairs). Table 10 in Appendix B provides an overview with additional statistics. We adapt the Faroese news sentiment dataset (Debess et al., 2024) (original labels: positive, negative, neutral) into minimal pairs by adding a sentiment-bearing sentence (“Hetta er gott/ringt”). Neutral labels are excluded as initial experiments showed that words such as “neutral” are never the most probable choice. We evaluate sentence- and article-level samples, keeping only items annotators of the original dataset agreed on, resulting in 91 sentence-level (55 positive, 36 negative) and 84 article-level (51 positive, 33 negative) pairs. Using the same dataset, we filter GPT-4-assigned topic labels confirmed by a human. Minimal pairs consist of one correct topic and one incorrect topic (not assigned to the article), with related-topic pairs curated to make the task realistic (e.g., Local News vs. International News). This yields 234 topic classification pairs. We also adapt the extractive QA dataset FoQA (Simonsen et al., 2025) (2,000 Faroese question–context–answer triplets) into minimal pairs via two methods: (1) Dataset Shuffling: Replacing the correct answer passage with an incorrect but plausible passage from another sample within the context, matching token length. This creates 21,867 pairs. (2) GPT-4 Adversarial Answers: Generating one alternative incorrect answer per sample that is also a span in the dataset, matching token length when possible. Exact length matches occurred for 611 answers; deviations were an average of 1.69 tokens longer than the correct answers. This yields 2,000 pairs.
@@ 3.4 Human Evaluation
Since no fine-grained evaluation sets exist for Faroese generation, we conduct a human evaluation to assess output quality across dimensions requiring subjective judgment: linguistic quality, naturalness, and contextual appropriateness. Given the resource-intensive nature of expert evaluation, we focus on four 360M models representing the best-performing configurations, varying along two dimensions: full fine-tuning versus LoRA, and transfer from Icelandic versus Mergeis+. Two native Faroese speakers serve as evaluators, both trained linguists with extensive NLP evaluation experience. We first followed a structured calibration protocol: evaluators jointly examined sample outputs, discussed edge cases, and iteratively refined the scoring guidelines until reaching consensus. Following this calibration phase, evaluators independently annotated a small set of approximately 10 samples, then met to resolve disagreements and clarify ambiguities before proceeding independently. Periodic check-ins were held throughout to maintain consistency without sharing individual scores.
Task Dimension $r$ $\kappa_{w}$ 95% CI Sent. Cont. Sem. coh. 0.427 0.383 [0.303, 0.460] Lex. cor. 0.409 0.286 [0.212, 0.357] Grammar 0.375 0.367 [0.277, 0.449] Fluency 0.298 0.295 [0.205, 0.379] Summ. Task compl. 0.728 0.718 [0.644, 0.781] Ling. qual. 0.829 0.819 [0.760, 0.866]
Table 2: Inter-annotator agreement: Pearson’s $r$ and quadratic-weighted Cohen’s $\kappa_{w}$ with bootstrapped 95% confidence intervals (10,00010{,}000 iterations). Sent. cont.: $n=400$, summ.: $n=200$.
@@
Given the limited capacity of small base models, we adopt sentence continuation as a straightforward generation task. Models are prompted with sentences drawn from a small manually compiled corpus of academic papers and local news articles (not included in FineWeb-2), with the final words removed and a trailing space added. All models produce running text, enabling direct comparison of output quality. A token cut-off of 100 was applied. Outputs are evaluated along four subdimensions of linguistic quality, each scored 0–5: lexical correctness (valid Faroese vocabulary, absence of hallucinated forms), grammatical accuracy (morphological and syntactic well-formedness, including spelling and typography), semantic coherence (meaningful, internally consistent content), and fluency/naturalness (native-like expression). Each annotator evaluated 400 continuations (100 prompts $\times$ 4 models). During guideline development, annotators agreed to rate the full generated output rather than the immediate completion alone, as the truncated sentence was typically resolved within a few words before the model continued generating further text. Inter-annotator agreement, measured by Pearson’s $r$, was 0.546 overall; all scores are reported in Table 2.
@@
Initial zero-shot experiments (i.e., prompting models directly to summarize a given text) failed to produce recognizable summaries, necessitating fine-tuning. As there was no Faroese summarization dataset, we constructed a synthetic one: 150 authentic texts spanning the same domains as the evaluation set (academic, news, blog), each paired with a summary generated by Claude Sonnet 4.444Fine-tuning details are provided in Table 8 in Appendix A. For evaluation, we selected 50 source texts not in FineWeb-2 and generated summaries from all four models (token cutoff: 400), yielding 200 summary–source pairs. Annotators rated two criteria: task completion (the extent to which the model successfully performed the summarization task) and linguistic quality, each rated on a 0–5 scale. Guidelines were developed through the same iterative calibration process as for sentence continuation, though the criteria were intentionally broader, focusing on overall task success and linguistic quality rather than fine-grained subdimensions. One point of disagreement during calibration concerned the threshold for assigning a score of 0, which required additional discussion to resolve. Inter-annotator agreement was substantially higher than for sentence continuation (Pearson’s r=0.879 overall; Table 2), likely because the coarser-grained criteria reduced room for subjective interpretation.
@@ 4.1 Benchmarks
We first give an overview of the benchmarks results with respect to their difficulty and reliability. FoBLiMP results are shown in Table 3. Models perform well on most linguistic acceptability probes, suggesting these tasks are relatively easy. The main exception is Translation Pairs, where scores are lower, possibly as translation error counts do not always reflect linguistic quality: translation errors can reflect aspects such as incorrect content compared to the source sentence.
Full LoRA 135M 360M 135M 360M En 95.25 96.55 96.12 96.55 +Da 96.98 95.25 98.27 96.12 +Is 97.41 99.13 97.41 98.27 +NoB 96.12 100.0 96.98 96.98 +NoN 96.12 98.27 95.68 97.41 +Sv 96.12 98.27 96.98 96.98 Mergeeq 97.41 98.27 96.55 95.25 Mergeis+ 97.84 97.84 97.41 96.55 Mergeda+is 97.84 97.41 97.41 99.13
@@
Full LoRA 135M 360M 135M 360M En 93.48 95.36 92.54 94.66 +Da 94.27 95.11 94.91 95.75 +Is 94.27 95.16 94.57 95.75 +NoB 95.06 95.31 94.91 95.55 +NoN 93.97 94.91 94.47 95.01 +Sv 94.61 95.85 95.36 95.80 Mergeeq 94.66 95.80 95.16 95.31 Mergeis+ 94.57 95.60 95.06 95.26 Mergeda+is 94.52 95.85 94.37 95.60
@@
(f) Mean of the individual FoBLiMP scores in 3(a)-3(e).
@@
Full LoRA 135M 360M 135M 360M En 60.43 68.13 60.43 72.52 +Da 72.52 75.82 61.53 76.92 +Is 63.73 75.82 60.43 69.23 +NoB 75.82 71.42 60.43 72.52 +NoN 76.92 74.72 67.03 71.42 +Sv 70.32 68.13 60.43 69.23 Mergeeq 75.82 72.52 60.43 62.63 Mergeis+ 68.13 80.21 60.43 63.73 Mergeda+is 76.92 75.82 62.63 65.93
@@
Full LoRA 135M 360M 135M 360M En 75.35 84.37 69.57 86.40 +Da 72.86 86.74 69.13 82.93 +Is 75.64 85.46 71.80 86.87 +NoB 72.04 84.16 70.00 86.60 +NoN 72.89 85.05 74.41 86.12 +Sv 70.82 85.12 70.42 86.10 Mergeeq 63.47 83.87 67.43 78.08 Mergeis+ 70.06 83.33 70.63 85.77 Mergeda+is 72.25 85.14 70.19 82.43
@@
(f) Mean of the individual FoBCoMP scores in 4(a)-4(e).
@@
Full LoRA 135M 360M 135M 360M En 4.98 3.75 5.51 4.48 +Da 4.19 3.56 4.25 3.55 +Is 4.44 3.48 4.53 3.53 +NoB 4.22 3.63 4.26 3.56 +NoN 4.60 3.66 4.90 4.08 +Sv 4.26 3.60 4.21 3.58 Mergeeq 4.08 3.41 4.61 3.93 Mergeis+ 4.08 3.41 4.58 3.77 Mergeda+is 4.22 3.49 4.56 3.80
Table 5: Average per-token perplexity on the Fineweb-2 evaluation set after CPT on Faroese.
@@ 4.2 RQ1: Transfer Languages are Important
Figure 1: Win rates of transfer languages across models and setups. In case of a tie, we count both.
@@
Models adapted via a Scandinavian transfer language consistently show lower perplexities than those adapted directly from English. The difference is larger for smaller models and for LoRA models, suggesting that LoRA benefits more from better-initialized parameters, consistent with Biderman et al. (2024) who find full tuning more sample-efficient than LoRA in domain adaptation. Table 5) shows that after adaptation on Faroese, Icelandic is the best individual transfer language for larger models, while for smaller models, higher-resource Mainland Scandinavian languages outperform Icelandic (except for Norwegian-Nynorsk, the lowest-resource Mainland Scandinavian language).
@@
Discussion
Icelandic emerges as the strongest transfer language overall. It yields the lowest perplexity prior to Faroese adaptation and shows clear advantages on FoBLiMP probes such as subject–verb agreement, supporting our decision to use Icelandic for human evaluation. Compared to English-only initialization, Icelandic provides almost consistent improvements on FoBLiMP (19/20 cases) and frequent gains on FoBCoMP (13/20 cases). We next examine how often all transfer setups outperform English, which would indicate that any additional transfer language is preferable to none. For FoBLiMP, transfer is particularly important for the 360M models: all transfer languages outperform English in 9/10 cases for 360M models, compared to 4/10 cases for 135M models. For FoBCoMP, however, the picture is less clear, with all transfer setups outperforming English in only 4/20 cases. When comparing all individual transfer languages (compare Figure 1), Icelandic achieves 14 wins overall and Danish 13. On FoBLiMP, Icelandic leads (9 wins vs. 7; 4 ties), whereas on FoBCoMP Danish performs better (9 wins vs. 6; 5 ties).
The overall finding that initializing Faroese models with a Scandinavian transfer language improves performance compared to English-only models confirms prior results for Faroese encoder models (Snæbjarnarson et al., 2023) and extends them to decoder models. Snæbjarnarson et al. (2023) report that an Icelandic BERT base model often outperforms a Danish base model, consistent with our results for perplexity and linguistic acceptability. Interestingly, they find that Danish BERT performs better on a semantic text similarity benchmark (the only semantic task in their work), aligning with our observation that Danish shows a slight tendency to outperform Icelandic on text comprehension tasks. More broadly, our findings are consistent with prior work on language transfer, which highlights both the importance of transfer languages and the benefits of using typologically related languages (Chronopoulou et al., 2023; Faisal and Anastasopoulos, 2022; Chang et al., 2024).